how does +3*x+9=0 ? I dont get it ....
Answer by Alan3354(69443)

how does +3*x+9=0 ? I dont get it ....
===============
If x = -3,
+3*(-3) + 9 = -9 + 9 = 0
------------------
The equation is true if x = -3
=================
Usually you see this:
3*x+9=0 solve for x
Subtract 9
3x = -9
Divide by 3
x = -3